Find all rational solutions of x4−6x3+22x2−30x+13=0


1
Expert's answer
2021-11-29T15:22:30-0500

"x^4\u22126x^3+22x^2\u221230x+13=0"

Let f(x) = x4−6x3+22x2−30x+13

"f(1) = 1-6+22-30+13=0"

So, "(x-1)" is a factor of "f(x)."



So, "x^4\u22126x^3+22x^2\u221230x+13 = (x-1)" ("x^3-5x^2+17x-13" )

Let "g(x)=x^3-5x^2+17x-13"

"g(1)=1-5+17-13=0"

So, "(x-1)" is a factor of "g(x)."



Thus, "f(x)=(x-1).g(x)=(x-1)(x-1)(x^2-4x+13)"

Now, put "(x^2-4x+13)=0"

Using Quadratic formula:

"x=\frac{4\pm\sqrt{16-52}}{2}=\frac{4\pm6i}{2}=2\pm3i"

Hence, the solution are "x=1,1,2+3i,2-3i" and rational solutions are "x=1,1"